# MetaQAP - A Meta-Learning Approach for Quality-Aware Pretraining in Image Quality Assessment

### v1

@@ 6 Conclusion
  • Funding: The funding for the study is provided by the Chinese Academy of Sciences President’s International Fellowship Initiative. Grant No.2024PVB0036.
@@
  • Code Availability: The code for the model and experiments is accessible via GitHub at:
    https://github.com/nisarahmedrana/MetaQAP.
  • Author Contributions: M.A.A. conceived and designed the study, developed the methodology, and led the writing of the manuscript. M.H. contributed to data analysis and experimental validation. N.A. developed the methodology and performed implementation, experiments and interpretation of results. G.S. assisted with dataset preparation, critical evaluation and pre-processing. Z.S. contributed to the development of the quality-aware loss function and supervised the pre-training experiments. H.H. and X.W. supported the implementation and fine-tuning of the CNN models. S.A. reviewed the statistical analyses and cross-dataset evaluations. W.J. provided oversight for the project, offered critical revisions to the manuscript, and ensured the research aligned with the broader objectives of the study. All authors reviewed and approved the final manuscript.

### v2

@@ 6 Conclusion
Acknowledgement
The authors extend their appreciation to the Deanship of Research and Graduate Studies at King Khalid University for funding this work through Large Group Project under grant number (RGP.2/556/45). Princess Nourah bint Abdulrahman University Researchers Supporting Project number (PNURSP2026R333), Princess Nourah bint Abdulrahman University, Riyadh, Saudi Arabia.
@@ Acknowledgement
  • Funding: The funding for the study is provided Deanship of Research and Graduate Studies at King Khalid University through Large Group Project under grant number (RGP.2/556/45) and Princess Nourah bint Abdulrahman University Researchers Supporting Project number (PNURSP2026R333).
@@
  • Code Availability: The code for the model and experiments is accessible via GitHub at: https://github.com/nisarahmedrana/MetaQAP.
  • Author Contributions: N.A. performed implementation, experiments, analysis of results and writeup of the manuscript. G.S. assisted with dataset preparation, critical evaluation and pre-processing. N. Alturki contributed to the development of the quality-aware loss function and supervised the pre-training experiments. N. Alasbali supported the implementation and fine-tuning of the CNN models. All authors reviewed and approved the final manuscript.